# Synthetic-Powered Predictive Inference

### v1

@@ Abstract
Conformal prediction is a framework for predictive inference with a distribution-free, finite-sample guarantee. However, it tends to provide uninformative prediction sets when calibration data are scarce. This paper introduces Synthetic-powered predictive inference (SPPI), a novel framework that incorporates synthetic data—e.g., from a generative model—to improve sample efficiency. At the core of our method is a score transporter: an empirical quantile mapping that aligns nonconformity scores from trusted, real data with those from synthetic data. By carefully integrating the score transporter into the calibration process, SPPI provably achieves finite-sample coverage guarantees without making any assumptions about the real and synthetic data distributions. When the score distributions are well aligned, SPPI yields substantially tighter and more informative prediction sets than standard conformal prediction. Experiments on image classification and tabular regression demonstrate notable improvements in predictive efficiency in data-scarce settings.
@@ 1.1 Background and motivation
At first glance, it may appear hard to use synthetic data to boost sample efficiency in a statistically valid way. After all, the distribution of synthetic data can be completely different from that of the data of interest. [^1] We refer to the limited dataset of interest as the real calibration set, to distinguish it from the (potentially synthetic) calibration data. Overcoming this challenge, we propose a principled framework that unlocks conformal prediction with the ability to incorporate synthetic data while preserving rigorous, model-agnostic, non-asymptotic coverage guarantees. Crucially, our method—SPPI—provides a coverage bound that requires no assumptions about the similarity between the real and synthetic data distributions. Still, when the distribution of synthetic and real scores is close, our approach yields a substantial boost in sample efficiency—resulting in more informative prediction sets than those produced by standard conformal prediction. A discussion of related literature is deferred to Appendix B.
@@ 1.2 Preview of the proposed method and our key contributions
Our key innovation is the introduction of the score transporter: a data-driven empirical quantile mapping function that transports the real calibration scores to resemble the synthetic scores. This mapping enables the construction of prediction sets for new test datapoints, leveraging the abundance of synthetic data. Crucially, the score transporter does not require data splitting, allowing full use of the real and synthetic calibration data. Furthermore, we develop a computationally efficient algorithm with a runtime complexity similar to that of standard conformal prediction. A pictorial illustration of our proposed calibration framework is provided in Figure S1.
We support our proposed SPPI framework with two theoretical guarantees. The first shows that when the synthetic and real score distributions are close, the achieved coverage closely matches the desired level. More generally, it characterizes how the distributional shift between the real and synthetic scores, and the construction of the score transporter affect the realized coverage.
@@
We demonstrate the practicality of our method on multi-class classification and regression tasks. For image classification on ImageNet, we explore two practical strategies for constructing synthetic data. The first leverages a generative model (Stable Diffusion [44] or FLUX [28]) to generate artificial images for each class. The second uses another set of real data, drawn from a different distribution, as the synthetic data. In the regression setting, we consider tabular panel data, using past panels as synthetic data and a recent panel as the real calibration set. Across all experiments, our method shows improvements in statistical efficiency—even when the real calibration set is very small, with as few as 15 datapoints. Software for reproducing the experiments is available at https://github.com/Meshiba/sppi.
@@ 3.1 Synthetic-powered predictive inference
In this section, we introduce our method—SPPI—which is designed to leverage the synthetic datapoints to effectively increase the sample size, thereby producing a non-conservative prediction set. We construct a split-conformal-type method that performs inference based on pre-constructed nonconformity scores. Throughout the section, we assume that the score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is fixed, and denote the real and synthetic scores as $S_{i}=s(X_{i},Y_{i})$ for $i\in[m+1]$ and $\smash{\tilde{S}_{j}=s(\tilde{X}_{j},\tilde{Y}_{j})}$ for $j\in[N]$, respectively. Here $S_{m+1}$ is the unobserved test score.
@@
where $\beta\in(0,1)$ is a predefined level, and $F:=F_{m,N,r}$ is defined as [^4] Here, for non-negative integers $a\leq b$, $\binom{b}{a}=b!/(a!(b-a)!)$ denotes the binomial coefficient, where $x!=x\cdot(x-1)\cdot\ldots\cdot 1$ is the factorial of a non-negative integer $x$.
@@
Roughly speaking, the score-transporter $T$ maps $\eta$ to a synthetic score in the corresponding window $I_{m}(r_{\eta})$ that is closest to $\eta$. The lower nearest neighbor $\mathrm{NN}_{m}^{-}$ is chosen carefully to act as a lower bound on the score, ensuring that the coverage can be tightly controlled. In Section D, we provide a detailed explanation of the motivation and intuition behind the score-transporter.
@@
Here $\tilde{Q}_{1-\alpha}$ is the $\lceil(1-\alpha)(N+1)\rceil$-th smallest score in $(\tilde{S}_{j})_{j\in[N]}$. We term this procedure Synthetic-powered predictive inference (SPPI). Building on the ideas of [55, 56], we extend our proposed method to achieve label-conditional coverage guarantees in Appendix F.
3.2 Simplifying the computation of SPPI
@@
Where $\tilde{Q}_{1-\alpha}^{\prime}$ is the $(\lceil{(1-\alpha)(N+1)}\rceil+1)$-th smallest score among $(\tilde{S}_{j})_{j\in[N]}$, and
@@
Based on this simplification, we present the complete SPPI procedure in Algorithm 1.
@@ 3.3 Theoretical guarantees
We now derive bounds on the coverage rate of the SPPI prediction set (9). The first bound shows that when the real and synthetic scores are similar (as measured by total variation distance), our method has a tight coverage around the desired level.
@@
Suppose the real calibration set $(X_{i},Y_{i})_{i\in[m]}$ is exchangeable with the test point $(X_{m+1},Y_{m+1})$ and the synthetic calibration set $(\tilde{X}_{j},\tilde{Y}_{j})_{j\in[N]}$ are drawn i.i.d., where the distribution $Q$ of their scores is continuous. Let $P_{(r)}^{m+1}$ and $Q_{(r)}^{m+1}$ denote the distribution of the $r$-th order statistic among $m+1$ i.i.d. draws from $P$ and $Q$, respectively. Then the prediction set $\smash{\widehat{C}(X_{m+1})}$ from (9) satisfies
@@
When the distributions $P$ and $Q$ differ greatly, the bounds in Theorem 3.3 may be loose, as they do not sufficiently account for the adjustment introduced by the map $T$ under distribution shift. Below, we provide alternative worst-case bounds for the coverage rate of the SPPI prediction set, which depend only on the sample sizes, and hold regardless of the relationship between $Q$ and $P$.
@@
While this result is somewhat non-explicit, the bounds can be computed fast, and remain close to the target level $1-\alpha$, as illustrated in Section D.1 through a set of plots. We emphasize that the bounds hold due to the careful construction of the score transport map from (8), and would not hold if we were to simply mix together the real and synthetic data. Moreover, the bounds imposes no condition on the distribution of the synthetic scores—it is even allowed for the synthetic scores to depend on the real calibration set. This provides significant flexibility in the choice of synthetic data even with a separate score function; see Section 3.4.
@@ 3.4 Improving the quality of synthetic scores
The quality of the SPPI prediction set depends on how well the distribution of the synthetic score $Q$ approximates the true score distribution $P$, as supported by Theorem 3.3. Thus, we can improve the prediction set by carefully constructing the synthetic scores. For example, we can seek a map $g$ such that the distribution of the adjusted synthetic score $\smash{\tilde{S}_{j}^{\prime}=g(\tilde{S}_{j})}$ better approximates the true distribution $P$.
@@ 3.4.2 Constructing a subset of synthetic data
Now, we shift to a different approach for improving the quality of synthetic scores: constructing a subset of the synthetic data that is more relevant for inference on the real data. This approach is particularly useful when the synthetic data comes from different sources, rather than sampled from a generative model. The idea is to select synthetic datapoints based on how well they approximate the real data. Then, we form a subset consisting of points with high approximation quality. Again, since Theorem 3.5 imposes no condition on the joint distribution of the synthetic scores, the bounds also hold for the SPPI prediction set constructed with this subset of synthetic scores.
@@ 4 Experiments
In this section, we compare the performance of the proposed SPPI procedure to that of standard conformal prediction in a setting where a small real calibration set and a large synthetic calibration set are available, each drawn from distinct and unknown distributions.
@@ Methods
We compare the following methods: OnlyReal—standard split conformal prediction [37] using the real calibration set; OnlySynth—conformal prediction applied to the synthetic calibration set as if it were real, which does not provide coverage guarantees; and SPPI (ours)—the proposed procedure outlined in Algorithm 1, applied with $\beta=0.4$.
@@ 4.1 Multi-class classification on the ImageNet data
4.1.1 SPPI with generated synthetic data
We use Stable Diffusion [44] to generate synthetic images resembling those in ImageNet. Examples and further details are provided in Figure S6 and Section H.4.
@@ 4.1.1 SPPI with generated synthetic data
In contrast, the proposed method, SPPI, achieves coverage within the theoretical bounds established in Theorem 3.5. For example, for the “Siberian husky” class, where the synthetic images differ significantly from the real ones, SPPI still produces informative prediction sets. For classes where the synthetic and real data are more aligned, such as the “lighter” class, SPPI shows low variance in coverage with smaller prediction set sizes.
We provide results for additional $\alpha$ levels in Section I.1. Further experiments regarding the effect of the size of the real calibration set on the performance of SPPI are shown in Section I.1.1. In addition, we provide experiments using the FLUX generative model [28], which exhibit similar trends to those observed with Stable Diffusion; see Section I.1.2. Examples of generated images and details are provided in Figures S7 and H.4, respectively.
Figure 1: Results for the ImageNet data: Coverage rates of OnlyReal, OnlySynth, and SPPI at target level $0.951-\alpha=$ 0.95, averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Table S3 for results across all classes.
4.1.2 SPPI with synthetic data from $k$-nearest subset selection
We now explore the performance of the subset-based variant of our approach, referred to as SPPI-Subset and described in Algorithm 2. The experiments in this section reflect scenarios where a generative model is unavailable.
@@ 4.1.2 SPPI with synthetic data from $k$-nearest subset selection
We apply the subset selection approach to improve the quality of the synthetic data, using a $k=20$ nearest-subset selection strategy, leading to $\tilde{N}=nk=300$ selected synthetic datapoints (see Algorithm 2). We compare this SPPI-Subset variant of our method to SPPI-Whole, where the latter denotes the SPPI procedure run with the entire synthetic set. Additionally, as a baseline, we include standard conformal prediction applied to the real set, OnlyReal.
@@
Both SPPI-Whole and SPPI-Subset achieve coverage within the theoretical bounds, generating smaller prediction sets compared to OnlyReal. In the label-conditional setting, SPPI-Subset achieves coverage that more closely aligns with the target $1-\alpha$ and produces smaller prediction sets, outperforming SPPI-Whole. This highlights the benefit of aligning the synthetic set more closely with the real distribution through $k$-nearest subset selection.
However, in the marginal setting, where the real calibration set includes images from 30 different classes, selecting a small subset of $k$ synthetic classes does not necessarily improve alignment with the real distribution. Consequently, SPPI-Subset, which uses only a subset of the synthetic data (300 images), exhibits higher variance in coverage compared to SPPI-Whole, which leverages the entire synthetic calibration set of 1,500 images.
Figure 2: Results for the ImageNet data: Coverage rates of OnlyReal, SPPI-Whole, and SPPI-Subset at target level $0.981-\alpha=$ 0.98, averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Table S8 for results across all classes.
We provide results for all classes appearing in the real data at additional values of the target level $\alpha$ in Section I.2. In Section I.2.1, we further illustrate the performance of the SPPI-Subset procedure for different values of the hyperparameter $k$.

### v2

@@ Abstract
Conformal prediction is a framework for predictive inference with a distribution-free, finite-sample guarantee. However, it tends to provide uninformative prediction sets when calibration data are scarce. This paper introduces Synthetic-powered predictive inference (SPI), a novel framework that incorporates synthetic data—e.g., from a generative model—to improve sample efficiency. At the core of our method is a score transporter: an empirical quantile mapping that aligns nonconformity scores from trusted, real data with those from synthetic data. By carefully integrating the score transporter into the calibration process, SPI provably achieves finite-sample coverage guarantees without making any assumptions about the real and synthetic data distributions. When the score distributions are well aligned, SPI yields substantially tighter and more informative prediction sets than standard conformal prediction. Experiments on image classification—augmenting data with synthetic diffusion-model generated images—and on tabular regression demonstrate notable improvements in predictive efficiency in data-scarce settings.
@@ 1.1 Background and motivation
At first glance, it may appear hard to use synthetic data to boost sample efficiency in a statistically valid way. After all, the distribution of synthetic data can be completely different from that of the data of interest. [^1] We refer to the limited dataset of interest as the real calibration set, to distinguish it from the (potentially synthetic) calibration data. Overcoming this challenge, we propose a principled framework that unlocks conformal prediction with the ability to incorporate synthetic data while preserving rigorous, model-agnostic, non-asymptotic coverage guarantees. Crucially, our method—SPI—provides a coverage bound that requires no assumptions about the similarity between the real and synthetic data distributions. Still, when the distribution of synthetic and real scores is close, our approach yields a substantial boost in sample efficiency—resulting in more informative prediction sets than those produced by standard conformal prediction. A discussion of related literature is deferred to Appendix B.
@@ 1.2 Preview of the proposed method and our key contributions
Our key innovation is the introduction of the score transporter: a data-driven empirical quantile mapping function that transports the real calibration scores to resemble the synthetic scores. This mapping enables the construction of prediction sets for new test datapoints, leveraging the abundance of synthetic data. Crucially, the score transporter does not require data splitting, allowing full use of the real and synthetic calibration data. Furthermore, we develop a computationally efficient algorithm with a runtime complexity similar to that of standard conformal prediction. A pictorial illustration of our proposed calibration framework is provided in Figure 1.
Figure 1: A high-level overview of the proposed method. The approach leverages a small labeled real dataset alongside a large labeled synthetic dataset. The score transporter maps scores from the real domain to the synthetic one. Calibration is then performed using the transported real scores and the synthetic scores.
We support our proposed SPI framework with two theoretical guarantees. The first shows that when the synthetic and real score distributions are close, the achieved coverage closely matches the desired level. More generally, it characterizes how the distributional shift between the real and synthetic scores, and the construction of the score transporter affect the realized coverage.
@@
We demonstrate the practicality of our method on multi-class classification and regression tasks. For image classification on ImageNet, we explore two practical strategies for constructing synthetic data. The first leverages a generative model (Stable Diffusion [45] or FLUX [29]) to generate artificial images for each class. The second uses another set of real data, drawn from a different distribution, as the synthetic data. In the regression setting, we consider tabular panel data, using past panels as synthetic data and a recent panel as the real calibration set. Across all experiments, our method shows improvements in statistical efficiency—even when the real calibration set is very small, with as few as 15 datapoints. Software for reproducing the experiments is available at https://github.com/Meshiba/spi.
@@ 3.1 Synthetic-powered predictive inference
In this section, we introduce our method—SPI—which is designed to leverage the synthetic datapoints to effectively increase the sample size, thereby producing a non-conservative prediction set. We construct a split-conformal-type method that performs inference based on pre-constructed nonconformity scores. Throughout the section, we assume that the score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is fixed, and denote the real and synthetic scores as $S_{i}=s(X_{i},Y_{i})$ for $i\in[m+1]$ and $\smash{\tilde{S}_{j}=s(\tilde{X}_{j},\tilde{Y}_{j})}$ for $j\in[N]$, respectively. Here $S_{m+1}$ is the unobserved test score.
@@
where $\beta\in(0,1)$ is a predefined level, and $F:=F_{m,N,r}$ is defined as [^4] Here, for non-negative integers $a\leq b$, $\binom{b}{a}=b!/(a!(b-a)!)$ denotes the binomial coefficient, where $x!=x\cdot(x-1)\cdot\ldots\cdot 1$ is the factorial of a non-negative integer $x$. Also, $p_{m,N,r}(k)$ is the probability mass function of the $r$-th order statistic from a random sample of size $m+1$ drawn without replacement from a finite population of size $N+m+1$ [e.g., 61, p. 243].
@@
Roughly speaking, the score-transporter $T$ maps $\eta$ to a synthetic score in the corresponding window $I_{m}(r_{\eta})$ that is closest to $\eta$. The lower nearest neighbor $\mathrm{NN}_{m}^{-}$ is chosen carefully to act as a lower bound on the score, ensuring that the coverage can be tightly controlled.
@@
Here $\tilde{Q}_{1-\alpha}$ is the $\lceil(1-\alpha)(N+1)\rceil$-th smallest score in $(\tilde{S}_{j})_{j\in[N]}$. We term this procedure Synthetic-powered predictive inference (SPI). Figure 2 presents a schematic overview of SPI with two candidate labels, illustrating each of the steps discussed above. Building on the ideas of [57, 58], we extend our proposed method to achieve label-conditional coverage guarantees in Appendix F.
(a) Candidate label $y_{1}$: the test score (purple) ranks second among the real scores. Its mapped synthetic neighbor—computed via (8) and outlined in purple—falls below the empirical quantile $\tilde{Q}_{1-\alpha}$, hence $y_{1}\in\widehat{C}(X_{m+1})$.
(b) Candidate label $y_{2}$: the test score (purple) ranks fourth among the real scores. Its mapped synthetic neighbor—computed via (8) and outlined in purple—exceeds the empirical quantile $\tilde{Q}_{1-\alpha}$, thus $y_{2}\notin\widehat{C}(X_{m+1})$.
Figure 2: Illustration of the synthetic-powered predictive inference for two candidate labels. Each panel displays sorted nonconformity scores: real scores on the left and synthetic scores on the right. The rectangle indicates the window in the synthetic space to which the test score can be mapped (as defined in (7)). The black-outlined circle indicates the $(1-\alpha)(1+\frac{1}{N})$th empirical quantile of the synthetic scores, $\tilde{Q}_{1-\alpha}$.
3.2 Simplifying the computation of SPI
@@
Here $\tilde{Q}_{1-\alpha}^{\prime}$ is the $(\lceil{(1-\alpha)(N+1)}\rceil+1)$-th smallest score among $(\tilde{S}_{j})_{j\in[N]}$, and
@@
Based on this simplification, we present the complete SPI procedure in Algorithm 1.
@@ 3.3 Theoretical guarantees
We now derive bounds on the coverage rate of the SPI prediction set (9). The first bound shows that when the real and synthetic scores are similar (as measured by total variation distance), our method has a tight coverage around the desired level.
@@
Suppose the real calibration set $(X_{i},Y_{i})_{i\in[m]}$ is exchangeable with the test point $(X_{m+1},Y_{m+1})$ and the synthetic calibration datapoints $(\tilde{X}_{j},\tilde{Y}_{j})_{j\in[N]}$ are drawn i.i.d., where the distribution $Q$ of their scores is continuous. Let $P_{(r)}^{m+1}$ and $Q_{(r)}^{m+1}$ denote the distribution of the $r$-th order statistic among $m+1$ i.i.d. draws from $P$ and $Q$, respectively. Then the prediction set $\smash{\widehat{C}(X_{m+1})}$ from (9) satisfies
@@
When the distributions $P$ and $Q$ differ greatly, the bounds in Theorem 3.3 may be loose, as they do not sufficiently account for the adjustment introduced by the map $T$ under distribution shift. Below, we provide alternative worst-case bounds for the coverage rate of the SPI prediction set, which depend only on the sample sizes, and hold regardless of the relationship between $Q$ and $P$.
@@
While this result is somewhat non-explicit, the bounds can be computed fast, and remain close to the target level $1-\alpha$, as illustrated in Section D.1 through a set of plots. We emphasize that the bounds hold due to the careful construction of the score transport map from (8), and would not hold if we were to simply mix together the real and synthetic data. Moreover, the bounds impose no condition on the distribution of the synthetic scores—it is even allowed for the synthetic scores to depend on the real calibration set. This provides significant flexibility in the choice of synthetic data even with a separate score function; see Section 3.4.
@@ 3.4 Improving the quality of synthetic scores
The quality of the SPI prediction set depends on how well the distribution of the synthetic score $Q$ approximates the true score distribution $P$, as supported by Theorem 3.3. Thus, we can improve the prediction set by carefully constructing the synthetic scores. For example, we can seek a map $g$ such that the distribution of the adjusted synthetic score $\smash{\tilde{S}_{j}^{\prime}=g(\tilde{S}_{j})}$ better approximates the true distribution $P$.
@@ 3.4.2 Constructing a subset of synthetic data
Now, we shift to a different approach for improving the quality of synthetic scores: constructing a subset of the synthetic data that is more relevant for inference on the real data. This approach is particularly useful when the synthetic data comes from different sources, rather than sampled from a generative model. The idea is to select synthetic datapoints based on how well they approximate the real data. Then, we form a subset consisting of points with high approximation quality. Again, since Theorem 3.5 imposes no condition on the joint distribution of the synthetic scores, the bounds also hold for the SPI prediction set constructed with this subset of synthetic scores.
@@ 4 Experiments
In this section, we compare the performance of the proposed SPI procedure to that of standard conformal prediction in a setting where a small real calibration set and a large synthetic calibration set are available, each drawn from distinct and unknown distributions.
@@ Methods
We compare the following methods: OnlyReal—standard split conformal prediction [38] using the real calibration set; OnlySynth—conformal prediction applied to the synthetic calibration set as if it were real, which does not provide coverage guarantees; and SPI (ours)—the proposed procedure outlined in Algorithm 1, applied with $\beta=0.4$.
@@ 4.1 Multi-class classification on the ImageNet data
4.1.1 SPI with generated synthetic data
We use Stable Diffusion [45] to generate synthetic images resembling those in ImageNet. Figure 3 shows representative examples, including images from an additional generative model discussed later. Additional examples and further details are provided in Figure S4 and Section H.4.
Figure 3: Examples of real and generated images for the golden retriever class. The first column displays a real ImageNet images, while the remaining columns show generated samples. The top row contains images generated by Stable Diffusion [45], and the bottom row by FLUX [29].
@@ 4.1.1 SPI with generated synthetic data
In contrast, the proposed method, SPI, achieves coverage within the theoretical bounds established in Theorem 3.5. For example, for the “Siberian husky” class, where the synthetic images differ significantly from the real ones, SPI still produces informative prediction sets. For classes where the synthetic and real data are more aligned, such as the “lighter” class, SPI shows low variance in coverage with smaller prediction set sizes.
We provide results for additional $\alpha$ levels in Section I.1. Further experiments regarding the effect of the size of the real calibration set on the performance of SPI are shown in Section I.1.1. In addition, we provide experiments using the FLUX generative model [29], which exhibit similar trends to those observed with Stable Diffusion; see Section I.1.2. Examples of generated images and details are provided in Figures 3, S5 and H.4, respectively.
Figure 4: Results for the ImageNet data: Coverage rates of OnlyReal, OnlySynth, and SPI at target level $0.951-\alpha=$ 0.95, averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Table S3 for results across all classes.
4.1.2 SPI with synthetic data from $k$-nearest subset selection
We now explore the performance of the subset-based variant of our approach, referred to as SPI-Subset and described in Algorithm 2. The experiments in this section reflect scenarios where a generative model is unavailable.
@@ 4.1.2 SPI with synthetic data from $k$-nearest subset selection
We apply the subset selection approach to improve the quality of the synthetic data, using a $k=20$ nearest-subset selection strategy, leading to $\tilde{N}=nk=300$ selected synthetic datapoints (see Algorithm 2). We compare this SPI-Subset variant of our method to SPI-Whole, where the latter denotes the SPI procedure run with the entire synthetic set. Additionally, as a baseline, we include standard conformal prediction applied to the real set, OnlyReal.
@@
Both SPI-Whole and SPI-Subset achieve coverage within the theoretical bounds, generating smaller prediction sets compared to OnlyReal. In the label-conditional setting, SPI-Subset achieves coverage that more closely aligns with the target $1-\alpha$ and produces smaller prediction sets, outperforming SPI-Whole. This highlights the benefit of aligning the synthetic set more closely with the real distribution through $k$-nearest subset selection.
However, in the marginal setting, where the real calibration set includes images from 30 different classes, selecting a small subset of $k$ synthetic classes does not necessarily improve alignment with the real distribution. Consequently, SPI-Subset, which uses only a subset of the synthetic data (300 images), exhibits higher variance in coverage compared to SPI-Whole, which leverages the entire synthetic calibration set of 1,500 images.
Figure 5: Results for the ImageNet data: Coverage rates of OnlyReal, SPI-Whole, and SPI-Subset at target level $0.981-\alpha=$ 0.98, averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Table S8 for results across all classes.
We provide results for all classes appearing in the real data at additional values of the target level $\alpha$ in Section I.2. In Section I.2.1, we further illustrate the performance of the SPI-Subset procedure for different values of the hyperparameter $k$.